$A$ die is thrown twice. What is the probability that
$(i)$ $5$ will not come up either time?
$(ii)$ $5$ will come up at least once?

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(A) Total number of outcomes $= 6 \times 6 = 36$.
$(i)$ Let $E$ be the event that $5$ comes up at least once. The outcomes where $5$ appears are $(5,1), (5,2), (5,3), (5,4), (5,5), (5,6), (1,5), (2,5), (3,5), (4,5), (6,5)$.
Number of favourable outcomes for $E = 11$.
Probability of $5$ coming up at least once,$P(E) = \frac{11}{36}$.
The probability that $5$ will not come up either time is $P(\text{not } E) = 1 - P(E) = 1 - \frac{11}{36} = \frac{25}{36}$.
$(ii)$ The probability that $5$ will come up at least once is $P(E) = \frac{11}{36}$.

Explore More

Similar Questions

If $P(\bar{B}) = 0.55$,then $P(B) = $ ...........

Two balanced dice are rolled simultaneously. Then,the probability that the numbers appearing on both the dice are different is ........

Difficult
View Solution

If the probability of an event is $p$,the probability of its complementary event will be

Two players,Sangeeta and Reshma,play a tennis match. It is known that the probability of Sangeeta winning the match is $0.62$. What is the probability of Reshma winning the match?

There are $6$ green,$5$ red,and $4$ blue identical balls in a bag. One ball is drawn at random from the bag. The probability that the ball drawn is not red is ...........

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo